# Quantum Brownian Motion as a Classical Stochastic Process in Phase Space

### v1

@@ I Introduction
Brownian motion—the erratic dynamics of a microscopic system interacting with a vast environment—is a cornerstone of statistical physics. Its classical description in terms of a Langevin equation Langevin1908 , combining friction and random noise, is ubiquitous across chemical kinetics, biophysics, and materials science Hanggi1990 . On smaller scales where quantum effects become significant, such as in low-temperature defect dynamics, quantum diffusion or tunneling Leggett1984QuantumTunneling ; BEC ; torres2022surface ; gottwald2015applicability ; Caldeira1983 , a fundamental question arises: how can Brownian motion be consistently quantized? This problem, known as quantum Brownian motion (QBM), is notoriously difficult. The principles of quantum mechanics forbid a naive quantization of classical dissipative equations, leaving a microscopic description of the system coupled to a macroscopic environment as the only rigorous path. However, the exponential growth of the Hilbert space dimension makes direct simulation intractable, creating a pressing need for a reduced description in terms of the Brownian particle alone.
The first microscopically consistent model for QBM was introduced by Caldeira and Leggett (CL) Caldeira1983path , representing the environment as a bath of independent harmonic oscillators linearly coupled to the particle. A foundational and well-known result of this model is that the Heisenberg equations of motion for the particle operators are formally identical to the classical generalized Langevin equation Weiss_2012 ; gardiner2004quantum . This formal correspondence suggests that the quantum dynamics might be representable in classical terms, yet translating this operator identity into a practical, all-temperature description of the quantum state has remained a challenge.
@@
The remainder of this paper is structured as follows: In Sec. II, we review the Caldeira-Leggett model and the exact path-integral framework. In Sec. III, we present the core result—the exact classical-stochastic mapping—and detail its proof. Section IV describes the numerical implementation. In Sec. V, we validate the method against known analytical results and contrast it with high-temperature master equations, highlighting the crucial all-temperature behavior. Section VI discusses the extension to non-quadratic potentials via the identified small parameter. We conclude in Sec. VII.
@@ II.1 Classical Caldeira-Leggett Model
The paradigmatic model for quantum Brownian motion is the Caldeira-Leggett (CL) model Caldeira1983path , which represents a one-dimensional Brownian particle of mass $m$ bilinearly coupled to a macroscopic environment. The environment is modeled as a bath of independent harmonic oscillators. The Hamiltonian for the total system is
@@ II.2.2 Reduced Quantum Dynamics with Intermediate Interventions
    Final State and Observables: After the last intervention at $t_{N}$, the state evolves to the final time $t_{f}$:
    $\rho_{SB}(t_{f})=U(t_{f}-t_{N})\,\rho_{SB}(t_{N}^{+})\,U^{\dagger}(t_{f}-t_{N}).$ (18)
    The physically relevant object is the reduced density matrix of the particle at $t_{f}$, obtained by tracing out the bath:
@@
    The expectation value of any particle observable $\hat{A}$ at an arbitrary time $t$ within the interval $0,t_{f}]$ is defined by the state at that time. The observable $\hat{A}$ acts solely on the particle’s Hilbert space. The time $t$ may precede, follow, or lie between the intervention times $\{t_{k}\}$. The general expression is
@@ III.2 The Reduced Stochastic Process for the Brownian Particle
$m\ddot{x}(t)=-\int_{-T}^{t}M(t-\tau)\dot{x}(\tau)d\tau+\xi^{(j)}(t),$ (25)
where the memory kernel $M(t)$ is defined in Eq. (10). The crucial difference from the classical case lies in the statistics of the stochastic force $\xi^{(j)}(t)$.
The force $\xi^{(j)}(t)$ is constructed from the sampled initial conditions of the bath oscillators $\{r_{i}^{(j)}(-T),p_{i}^{(j)}(-T)\}$:
$\xi^{(j)}(t)=\sum_{i}c_{i}\Big[$ $r_{i}^{(j)}(-T)\cos(\omega_{i}(t+T))$ (26)
@@
Because these initial conditions are drawn from the quantum thermal Wigner functions $W_{i}^{\beta}$, the ensemble-averaged noise correlation is given by the quantum fluctuation-dissipation theorem:
$\langle\xi(t)\xi(s)\rangle_{\text{ens}}=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\\ \times\left[\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)]-i\sin[\omega(t-s)]\right].$ (26)
The symmetric (real) part of this correlation, which governs observable averages, is
$\frac{1}{2}\langle\{\xi(t),\xi(s)\}\rangle_{\text{ens}}=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)].$ (27)
@@ III.3 Incorporating State Preparation and Interventions
The preparation formalism of Section 2.2 is incorporated naturally within this trajectory picture. A preparation or intervention at time $t_{k}$, described by operators $\{O_{\alpha}^{(k)}\}$, corresponds to a non-classical update of the particle’s Wigner function. In our stochastic framework, this is implemented by assigning a complex weight $w_{\alpha}^{(j)}$ to each trajectory $j$ at time $t_{k}$.
The weight is derived from the preparation function $\lambda$ in the Wigner representation. For a trajectory that has phase-space coordinates $(\bar{r},\bar{p})$ just before $t_{k}$ and $(r_{0},p_{0})$ just after, the weight contributed by the intervention is
$w^{(j)}(t_{k})=\lambda(r_{0},p_{0}|\bar{r},\bar{p}).$ (29)
@@ IV.1 Generation of Quantum Noise
$z^{*}(\omega)=z(-\omega),\quad\langle z(\omega)z^{*}(\omega^{\prime})\rangle=\delta(\omega-\omega^{\prime}),\quad\langle z(\omega)z(\omega^{\prime})\rangle=0.$ (33)
@@ V.1 Benchmark: Gaussian State Preparation
As a first test, we consider a Gaussian preparation that localizes the particle’s position Grabert1988 . In coordinate representation, the preparation function is
$\lambda(r_{0},\bar{r},q_{0},\bar{q})=\delta(q_{0}-\bar{q})\delta(r_{0}-\bar{r})\frac{1}{\sqrt{2\pi\sigma_{0}^{2}}}\exp\left(-\frac{r_{0}^{2}}{2\sigma_{0}^{2}}-\frac{q_{0}^{2}}{8\sigma_{0}^{2}}\right),$ (39)
which corresponds to a state centered at $\langle x\rangle=0$ with initial spatial variance $\sigma_{0}^{2}$. The uncertainty principle mandates a corresponding increase in momentum variance, reflected in the Wigner representation:
$\lambda(r_{0},\bar{r},p_{0},\bar{p})=\delta(r_{0}-\bar{r})\frac{1}{2(2\pi)^{2}}\exp\left(-\frac{r_{0}^{2}}{2\sigma_{0}^{2}}-\frac{2\sigma_{0}^{2}}{\hbar^{2}}(p_{0}-\bar{p})^{2}\right).$ (40)
@@
Figure 1: Time evolution of the coordinate dispersion $\sigma^{2}(t)$ for a Gaussian preparation at $T=0$. Points: stochastic Monte Carlo simulation. Solid line: analytical result from Eq. (41). The agreement validates the numerical method. Parameters: $\sigma_{0}=1$, $\gamma=\pi/2$, $\varepsilon=0.5$, $m=\hbar=1$.
@@ V.2 Decoherence Dynamics: Schrödinger Cat State
$\begin{gathered}|\psi\rangle=\frac{1}{\sqrt{\mathcal{N}}}(|+\rangle+|-\rangle),\\ \langle x|\pm\rangle=\frac{1}{(2\pi\sigma^{2})^{1/4}}\exp\left(-\frac{(x\pm x_{0})^{2}}{4\sigma^{2}}\right),\end{gathered}$ (42)
@@
Figure 2: Decoherence of a Schrödinger cat state. The observable $\langle O\rangle$ (Eq. (43)) decays fastest for the exact quantum noise (solid line). The high-temperature master equation (dashed) and white-noise process (dotted) give similar, slower decoherence, failing to capture the enhanced quantum fluctuations at low temperature. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
@@ V.3 Low-Temperature Thermalization and the Role of the Cutoff
Figure 3: Momentum dispersion $\langle p^{2}(t)\rangle$ relaxation for Schrödinger cat state. The exact quantum result (top lines, analytical and stochastic) shows rapid initial growth on a timescale $\sim\varepsilon$, absent in the high-temperature white-noise model (dashed). This demonstrates the quantum bath’s role in fast early-time dynamics. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
@@ A.2 Quantum influence functional with factorized initial condition
$K_{Q}(r_{f}p_{f}t_{f}|r_{i}p_{i}t_{i})=\int dq_{i}dq_{f}e^{-\frac{i}{\hbar}(p_{f}q_{f}-p_{i}q_{i})}\int_{r_{i}}^{r_{f}}Dr\int_{q_{i}}^{q_{f}}Dq\\ \exp[-\frac{im}{\hbar}\int_{t_{i}}^{t_{f}}dsq(s)\bigg(\ddot{r}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{r}(u)\bigg)\\ -\frac{1}{2\hbar^{2}}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duq(s)N(s-u)q(u)\bigg]e^{\frac{im}{\hbar}(\dot{r}_{f}q_{f}-\dot{r}_{i}q_{i})}$ (61)

### v2

@@ I Introduction
Brownian motion—the erratic dynamics of a microscopic system interacting with a vast environment—is a cornerstone of statistical physics. Its classical description in terms of a Langevin equationLangevin1908 , combining friction and random noise, is ubiquitous across chemical kinetics, biophysics, and materials science Hanggi1990 . On smaller scales where quantum effects become significant, such as in low-temperature defect dynamics, quantum diffusion or tunneling Leggett1984QuantumTunneling ; BEC ; torres2022surface ; gottwald2015applicability ; Caldeira, a fundamental question arises: how can Brownian motion be consistently quantized? This problem, known as quantum Brownian motion (QBM), is notoriously difficult. The principles of quantum mechanics forbid a naive quantization of classical dissipative equations, leaving a microscopic description of the system coupled to a macroscopic environment as the only rigorous path. However, the exponential growth of the Hilbert space dimension makes direct simulation intractable, creating a pressing need for a reduced description in terms of the Brownian particle alone.
The first microscopically consistent model for QBM was introduced by Caldeira and Leggett (CL) Caldeira1983 , representing the environment as a bath of independent harmonic oscillators linearly coupled to the particle. A foundational and well-known result of this model is that the Heisenberg equations of motion for the particle operators are formally identical to the classical generalized Langevin equation Weiss_2012 ; gardiner2004quantum . This formal correspondence suggests that the quantum dynamics might be representable in classical terms, yet translating this operator identity into a practical, all-temperature description of the quantum state has remained a challenge.
@@
The remainder of this paper is structured as follows: In Sec. II, we review the Caldeira-Leggett model. In Sec. III, we present the core result—the exact classical-stochastic mapping—and detail its proof. Section IV describes the numerical implementation. In Sec. V, we validate the method against known analytical results and contrast it with high-temperature master equations, highlighting the crucial all-temperature behavior. Section VI discusses the extension to non-quadratic potentials via the identified small parameter. We conclude in Sec. VII.
@@ II.1 Classical Caldeira-Leggett Model
The paradigmatic model for quantum Brownian motion is the Caldeira-Leggett (CL) model Caldeira1983 , which represents a one-dimensional Brownian particle of mass $m$ bilinearly coupled to a macroscopic environment. The environment is modeled as a bath of independent harmonic oscillators. The Hamiltonian for the total system is
@@ II.2.1 Joint State and Preparation
In coordinate representation it is described by preparation function
$\lambda(x,x^{\prime}|\bar{x},\bar{x}^{\prime})=\sum_{j}\langle x^{\prime}|O_{j}|\bar{x}^{\prime}\rangle\langle\bar{x}|O_{j}^{\dagger}|x\rangle.$ (15)
@@ II.2.2 Reduced Quantum Dynamics with Intermediate Interventions
    Final State and Observables: After the last intervention at $t_{N}$, the state evolves to the any time $t>t_{N}$:
    $\rho_{SB}(t)=U(t-t_{N})\,\rho_{SB}(t_{N}^{+})\,U^{\dagger}(t-t_{N}).$ (19)
    The physically relevant object is the reduced density matrix of the particle at $t$, obtained by tracing out the bath:
@@
    The expectation value of any particle observable $\hat{A}$ at an arbitrary time $t$ within the interval $[0,t]$ is defined by the state at that time. The observable $\hat{A}$ acts solely on the particle’s Hilbert space. The time $t$ may precede, follow, or lie between the intervention times $\{t_{k}\}$. The general expression is
@@ III.2 The Reduced Stochastic Process for the Brownian Particle
$m\ddot{x}(t)=-\int_{-T}^{t}M(t-\tau)\dot{x}(\tau)d\tau+\xi_{W}^{(j)}(t),$ (26)
where the memory kernel $M(t)$ is defined in Eq. (10). The crucial difference from the classical case lies in the statistics of the stochastic force $\xi_{W}^{(j)}(t)$.
The force $\xi_{W}^{(j)}(t)$ is constructed from the sampled initial conditions of the bath oscillators $\{r_{i}^{(j)}(-T),p_{i}^{(j)}(-T)\}$:
$\xi_{W}^{(j)}(t)=\sum_{i}c_{i}\Big[$ $r_{i}^{(j)}(-T)\cos(\omega_{i}(t+T))$ (27)
@@
Because these initial conditions are drawn from the quantum thermal Wigner functions $W_{i}^{\beta}$, the ensemble-averaged noise correlation is given by the quantum fluctuation-dissipation theorem for noise operator:
$\langle\hat{\xi}(t)\hat{\xi}(s)\rangle=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\\ \times\left[\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)]-i\sin[\omega(t-s)]\right].$ (28)
The symmetric (real) part of this correlation, which governs observable averages, for sampled $\xi_{W}$ is
$\langle\xi_{W}(t)\xi_{W}(s)\rangle=\frac{1}{2}\langle\{\hat{\xi}(t),\hat{\xi}(s)\}\rangle=\\ =\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)].$ (29)
@@ III.3 Incorporating State Preparation and Interventions
The preparation formalism of Section 2.2 is incorporated naturally within this trajectory picture. A preparation or intervention at time $t_{k}$, described by operators $\{O_{\alpha}^{(k)}\}$, corresponds to a non-classical update of the particle’s Wigner function. In our stochastic framework, this is implemented by assigning a weight $w_{\alpha}^{(j)}$ to each trajectory $j$ at time $t_{k}$.
The weight is derived from the preparation function $\lambda$ in the Wigner representation Hanggi1993 . For a trajectory that has phase-space coordinates $(\bar{r},\bar{p})$ just before $t_{k}$ and $(r_{0},p_{0})$ just after, the weight contributed by the intervention is
$w^{(j)}(t_{k})=\lambda(r_{0}^{(j)},p_{0}^{(j)}|\bar{r}^{(j)},\bar{p}^{(j)}),$ (31)
where $\lambda(r_{0},p_{0}|\bar{r},\bar{p})=\int d\bar{q}dq_{0}e^{\frac{i}{\hbar}(\bar{q}\bar{p}-q_{0}p_{0})}\lambda(r_{0},q_{0}|\bar{r},\bar{q})$. The point $(r_{0},p_{0})$ is assumed to be uniformly distributed in phase space, and the conditional probability of a trajectory break between points $(\bar{r},\bar{p})$ and $(r_{0},p_{0})$ is taken into account in the weight. This allows for working with essentially quantum preparation functions that have negative regions and are not interpreted as probabilities.
@@ IV.1 Generation of Quantum Noise
$z^{*}(\omega)=z(-\omega),\quad\langle z(\omega)z^{*}(\omega^{\prime})\rangle=\delta(\omega-\omega^{\prime}),\\ \langle z(\omega)z(\omega^{\prime})\rangle=0.$ (35)
@@ V.1 Benchmark: Gaussian State Preparation
As a first test, we consider a Gaussian preparation that localizes the particle’s position which is centered at $\langle x\rangle=0$ with initial spatial variance $\sigma_{0}^{2}$ Grabert1988 . The uncertainty principle mandates a corresponding increase in momentum variance, reflected in the Wigner representation:
$\lambda(r_{0},\bar{r},p_{0},\bar{p})=\frac{\delta(r_{0}-\bar{r})}{2(2\pi)^{2}}\exp\left(-\frac{r_{0}^{2}}{2\sigma_{0}^{2}}-\frac{2\sigma_{0}^{2}}{\hbar^{2}}(p_{0}-\bar{p})^{2}\right).$ (41)
@@
Figure 1: Time evolution of the coordinate dispersion $\sigma^{2}(t)$ for a Gaussian preparation at $T=0$. Blue line: stochastic Monte Carlo simulation. Orange line: analytical result from Eq. (42). The agreement validates the numerical method. Parameters: $\sigma_{0}=1$, $\gamma=\pi/2$, $\varepsilon=0.5$, $m=\hbar=1$.
@@ V.2 Decoherence Dynamics: Schrödinger Cat State
$\begin{gathered}|Cat\rangle=\frac{1}{\sqrt{\mathcal{N}}}(|+\rangle+|-\rangle),\\ \langle x|\pm\rangle=\frac{1}{(2\pi\sigma^{2})^{1/4}}\exp\left(-\frac{(x\pm x_{0})^{2}}{4\sigma^{2}}\right),\end{gathered}$ (43)
@@
Figure 2: Decoherence of a Schrödinger cat state. The observable $\langle O\rangle$ (Eq. (44)) decays fastest for the exact quantum noise (blue line). The high-temperature master equation (orange) and white-noise process (green) give similar, slower decoherence, failing to capture the enhanced quantum fluctuations at low temperature. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
@@ V.3 Low-Temperature Thermalization and the Role of the Cutoff
Figure 3: Momentum dispersion $\langle p^{2}(t)\rangle$ relaxation for Schrödinger cat state. The exact quantum result (blue and orange lines, analytical and stochastic) shows rapid initial growth on a timescale $\sim\varepsilon$, absent in the high-temperature white-noise model (green). This demonstrates the quantum bath’s role in fast early-time dynamics. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
@@ Appendix A Influence functional
In this section, we show the equality of the GLE approach and the exact quantum solution using the influence functional Grabert1988 .
@@ A.2 Quantum influence functional with factorized initial condition
$K_{Q}(r_{f}p_{f}t_{f}|r_{i}p_{i}t_{i})=\int dq_{i}dq_{f}e^{-\frac{i}{\hbar}(p_{f}q_{f}-p_{i}q_{i})}\int_{r_{i}}^{r_{f}}Dr\\ \int_{q_{i}}^{q_{f}}Dq\exp[-\frac{im}{\hbar}\int_{t_{i}}^{t_{f}}dsq(s)\bigg(\ddot{r}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{r}(u)\bigg)\\ -\frac{1}{2\hbar^{2}}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duq(s)N(s-u)q(u)\bigg]e^{\frac{im}{\hbar}(\dot{r}_{f}q_{f}-\dot{r}_{i}q_{i})}$ (62)